During strenuous exercise an athlete can burn 12 calories per minute on the treadmill and 8.5 calories per minute on the elliptical. The athlete uses both machines and burns 325 calories in 30 minutes. How many minutes does the athlete spend on each machine?
Found 2 solutions by fractalier, Cromlix:
Answer by fractalier(6550)

Call the minutes spent on the treadmill and elliptical, t and e. Then we write
t + e = 30
12t + 8.5e = 325
Now multiply the top one by 12 and subtract from the bottom one...we get
12t + 8.5e = 325
-(12t + 12e = 360)
-------------------
-3.5e = -35
e = 10 minutes on the elliptical which means
t = 20 minutes on the treadmill

Answer by Cromlix(4381)

Hi there,
Make treadmill = x
and elliptical = y
x + y = 30............(1)
12x + 8.5y = 325......(2)
Multiply Eq(1) by 8.5
8.5x + 8.5y = 255.....(1)
12x + 8.5y = 325......(20
Subtract Eq(1) from Eq(2)
3.5x = 70
x = 20
Substitute x = 20 into
Eq(1)
x + y = 30
20 + y = 30
y = 30 - 20
y = 10
The athlete spends 20 minutes
on the treadmill and 10 minutes
on the elliptical.
Hope this helps :-)